Robert Johnson owns a farm and breeds race horses. To purchase three thoroughbreds,
he signs a 180-day note with a maturity value of $ 255,000 and proceeds of $ 243,137.50.
Find (a) the discount and (b) the true (or effective) rate .
1
Expert's answer
2020-10-12T17:18:26-0400

(a)

The discount is "3\cdot(255000-243137.50)=35587.50"  ($).


(b)

180 days is 6 month, the nominal rate is "\frac{255000-243137.50}{243137.50}≈0.04879=4.879\%",

so the true (or effective) rate is "(1+\frac{0.04879}{6})^6-1≈0.04979=4.979\%".